In this question you will find the steady-state probability distribution for the regular transistion matrix below with 3 states A, B, and C.

  A      B      C
A
B
C
0.3 0.3 0.4
0.5 0.0 0.5
0.4 0.2 0.4

Give the following answers as fractions OR as decimals correct to at least 5 decimal places.

What is the long term probability of being in state A?


What is the long term probability of being in state B?


What is the long term probability of being in state C?
Answer #1
let fixed probability vector is W = [a b c]
therefore from relation: WT =W
0.3a+0.5b+0.4c=a…..(1)
0.3a+0b+0.2c=b…..(2)
0.4a+0.5b+0.4c=c…..(3)
solving above: a=0.38168 ; b=0.19847 ; c=0.41985

  long term probability of being in state A =0.38168

long term probability of being in state B =0.19847

long term probability of being in state C=0.41985